OpenStudy (anonymous):

What is the surface area of a rectangular prism with length 13mm, width 2mm, and height 7mm? 44mm(to the 2nd power) 132mm(to the 2nd power) 182(to the 2nd power) 262mm(to the 2nd power)

OpenStudy (anonymous):

is there any picture of the dimensions rough diagram

OpenStudy (anonymous):

using the formula SA = 2bh + 2bw + 2hw b = 13 w = 2 h = 7 SA = 2(13)(7) + 2(13)(2) + 2(7)(2) SA = 182 + 52 + 28 SA = 262 Then we add the mm since that is the one were using and the ^2 because it is a surface area. Answer 262 mm^2

OpenStudy (anonymous):

Thank you so much ^-^ @jennyfer